### 09 Mar 2026 v4

@@ Abstract
Black Holes (BHs) with integrable singularities (IS) have attracted attention because, unlike usual regular BHs, they avoid a potentially unstable de Sitter core and the presence of an internal horizon that breaks predictability, while exhibiting finite tidal forces that allow nondestructive radial infall. On the other hand, the idea that an interior region, rather than a pointlike mass, can generate a Schwarzschild exterior BH region has recently gained attention Ovalle (2024); Maeda (2025). In this work, motivated by the existence of an inner horizon and a non integrable singularity in the Reissner Nordström solution (RN), we address the possibility that the RN BH possesses an interior region with an IS and no inner horizon, instead of a pointlike mass. We propose that the nature of this interior region is given by a cloud of strings (CS) with IS and by a new fluid of strings (FS) with IS, determining the physical properties that both must satisfy in order to support such a configuration. To this end, we establish the matching conditions between these interior regions with IS and the RN exterior BH solution. This leads to the requirement that the continuity of temperature must be satisfied at the interface, while discontinuities in the tangential pressure signal phase transitions. Thus, for both interior regions, we determine the conditions relating them to the RN exterior region that must be satisfied in order to physically represent the interior region with an IS while fulfilling the matching conditions. Furthermore, we identify the conditions under which phase transitions occur or do not occur.
@@ I Introduction
The observation of gravitational waves produced by the merger of two spinning black holes Abbott et al. (2016) has firmly positioned black holes among the most compelling phenomena in gravitational physics. In general, black hole solutions are characterized by the presence of a central singularity, where the spacetime description breaks down, signaled by the divergence of the curvature invariants. To address this problem, the standard approach involves specifying a particular form of matter within the energy-momentum tensor, which gives rise to regular black holes (RBHs). It is well established that RBH solutions typically include an inner horizon, which can also lead to the presence of a de Sitter–like core near the origin. In this context, several studies have indicated that the inner horizon is inherently unstable Poisson and Israel (1990); Brown et al. (2011). Along the same lines, a more recent study, reference Carballo-Rubio et al. (2023), argues that instabilities at the cores of regular black holes associated with the inner horizon are inevitable, since the mass inflation phenomenon represents a fundamental feature of RBHs with astrophysical significance. On the other hand, it is also worth noting that, as claimed in reference Ovalle (2023), predictability breaks down after crossing the inner horizon.
As an alternative to RBH solutions, Ref. Lukash and Strokov (2013) proposes a gravitational scenario in which a particle can reach the radial origin without being destroyed. Rather than a potentially unstable de Sitter core associated with an inner horizon, this model exhibits an integrable singularity at the origin. Integrable singularities are characterized by a divergence of the Ricci scalar while their volume integrals remain finite near the center. The elimination of the inner horizon thus prevents possible instabilities. Furthermore, recent studies in Ref. Estrada et al. (2024) show that a central integrable singularity allows for finite tidal forces, extendable radial geodesics, and a weak singularity, as formalized by Tipler’s extension theorem Nolan (2000). This ensures that an object approaching the central singularity would not undergo spaghettification. In addition, Ref. Casadio et al. (2023) offers a quantum perspective on the existence of such integrable singularities. For other recent applications, see Refs. Ovalle (2023); Estrada (2025); Arrechea et al. (2025).
On the other hand, String theory posits that the fundamental constituents of the universe are one-dimensional objects called strings Younesizadeh and Jokar (2022). The dynamics of a single 1-brane string can be extended to an ensemble Yang et al. (2023), forming a cloud of strings (CS). Embedding such a cloud of strings modifies the spacetime geometry, producing black hole solutions surrounded by cloud of strings Letelier (1979), and several recent studies have analyzed black holes interacting with string clouds Zafar et al. (2025); Javed et al. (2026); Gogoi et al. (2025); Ahmed et al. (2025); Cai et al. (2025). A generalization of this model includes pressure or tension in the energy–momentum tensor, giving it a structure similar to a perfect fluid, i.e., a fluid of strings (FS) Letelier (1983). In this connection, recently an equation of state $\rho(r)=\alpha(r)p(r)$ was proposed Nunes dos Santos (2025) for a fluid of strings, leading to a regular black hole model.
Following the definition of the conserved energy associated with a timelike Killing vector given in Ref. Aoki et al. (2021), we note that this conserved energy diverges for the standard cloud of strings energy density. Motivated by this, we first introduce in this work a modified version of the string-cloud energy density within the fluid-of-strings framework. This modification can be interpreted as a geometrical screening of the original distribution. As a result, this leads to the convergence of the conserved energy charge. In this way, we first provide a new fluid-of-strings black hole model, constructed through the equation of state $\rho(r)=\alpha(r)p(r)$. Furthermore, we show that the Ricci scalar near the origin behaves in a way consistent with an integrable singularity, and that the new FS model is also free of a potentially unstable inner horizon.
On the other hand, the Schwarzschild solution, which according to Birkhoff’s theorem Hawking and Ellis (2023) is the unique spherically symmetric vacuum solution of Einstein’s equations, possesses a strong central singularity. As pointed out in reference Ovalle (2024), the Schwarzschild black hole features an event horizon, which can be viewed as a membrane of radius $h$ separating two causally disconnected space-time regions: an interior and an exterior one. The interior region, defined by $r<h$, contains the central singularity shielded by the event horizon. In this domain, the metric function $f(r)$ becomes negative, causing the radial and temporal coordinates to interchange their roles and thereby revealing the dynamical character of the interior geometry. As a consequence, gravitational collapse toward the central singularity becomes unavoidable, in agreement with Penrose’s theorem Penrose (1965). Conversely, an observer in the exterior region has no causal access to the interior of the black hole: once the boundary at $r=h$ is crossed, it is impossible to return or communicate with the exterior. In this line of research, remarkably, the idea that an interior region, rather than a pointlike mass, can generate a Schwarzschild black hole exterior has recently gained attention Ovalle (2024); Maeda (2025). This configuration is obtained by matching an interior solution to the Schwarzschild exterior at the event horizon, without introducing a lightlike thin shell. Reference Ovalle (2024) examines whether an alternative structure, beyond a pointlike mass, can give rise to the Schwarzschild black hole region for $r\geq h$. In particular, the authors aim to describe an interior configuration associated with this region while keeping the exterior Schwarzschild geometry unchanged and ensuring that tidal forces remain finite everywhere. In connection with the latter, the authors propose that the internal region may contain either a central singularity or a regular core. Reference Maeda (2025) constructs a so-called fake Schwarzschild model, observationally indistinguishable from the Schwarzschild black hole, by matching a dynamical spherically symmetric interior containing matter to the Schwarzschild exterior at the event horizon without introducing a lightlike thin shell. It is shown that this configuration can be obtained from a perfect-fluid solution satisfying a linear equation of state of the type considered in Ref. Maeda (2023), analytically continued inside the horizon.
Reissner Nordström solution represents an electrically charged BH, arising from electromagnetic sources in the energy momentum tensor and characterized by the presence of an inner horizon and a central singularity. Motivated by the considerations described in this introduction, and by the existence of an inner horizon and a non-integrable singularity in the Reissner–Nordström solution (RN), in this work, we study the physical and thermodynamic properties of a scenario in which the RN BH possesses an interior region with an integrable singularity and no inner horizon, instead of a pointlike mass. We investigate the possibility that the nature of the matter sources capable of supporting such an internal geometry with an integrable singularity corresponds to those associated with clouds of strings and with the new string fluid introduced in this work.
To carry out this work, we will first define the conditions that a general internal geometry must satisfy in order to represent an integrable singularity, as well as the Israel Darmois Israel (1966) junction conditions between a general internal region and a general external BH geometry. We will then examine the connection between the junction conditions and the thermodynamic properties of BHs, and explore their relation to the existence of phase transitions. In this context, we will show that satisfying the junction conditions leads to the requirement that the continuity of temperature must hold at the interface, while discontinuities in the tangential pressure signal phase transitions. It is worth mentioning that these general conditions could also be of physical interest for future studies of other BH exterior regions, instead of RN, that possess a central singularity, for example BHs with cosmic void density profiles Lustosa et al. (2025), BHs surrounded by dark matter Xu et al. (2018), hairy BHs Ovalle et al. (2021), and quintessential BHs Kiselev (2003); Estrada et al. (2025). In this way, applying the general conditions described above, in this work we focus on the physical and thermodynamic properties of the mentioned configuration where the exterior geometry is given by the Reissner Nordström BH, while the interior geometry is described by a cloud and by a new fluid of strings with integrable singularities. Specifically, we will determine the physical properties that both the cloud of strings (CS) and the new fluid of strings (FS) must satisfy in order to represent the interior of an exterior RN BH region. Using the general analysis described above, we will determine the relations between the parameter spaces of both interior regions and the RN exterior BH for which the matching conditions are satisfied and, as a consequence, the temperature remains continuous at the interface. In addition, for both interior regions we will determine the conditions relating them to the RN exterior region that must be satisfied for phase transitions to occur or not occur.
@@
II A new Fluid of strings black hole model with an integrable singularity
@@ III.2 Description of the exterior spacetime
IV INTERIOR REGION AS AN INTEGRABLE SINGULARITY
@@
V Junction conditions
@@
VI Example 1: A Reissner–Nordström Exterior with a Cloud of Strings as the Interior Region with an Integrable Singularity
@@ VI.2 Junction conditions:
VII Example 2: A Reissner–Nordström Exterior with a new fluid of Strings as the Interior Region with an Integrable Singularity
@@ VIII Discussion and Summary
In this work, motivated by the existence of an inner horizon and a non-integrable singularity in the Reissner–Nordström solution (RN), we study the physical and thermodynamic properties of a scenario in which the RN black hole possesses an interior region with an integrable singularity and no inner horizon, instead of a pointlike mass. We have investigated the possibility that the nature of the matter sources capable of supporting such an internal geometry with an integrable singularity (IS) corresponds to clouds of strings (CS) and fluids of strings (FS). In this connection, motivated by the fact that the conserved charge energy in the CS case diverges, we have proposed a model for the internal region based on an FS in which the energy density, which can be interpreted as a geometrical screening of the original CS distribution, leads to a finite conserved charge energy. We have shown that both the CS and FS models used in this work can represent an interior region with an IS and that they lack an inner horizon where predictability would be lost.
To carry out this work, in Section IV we have determined the conditions that a general internal geometry must satisfy in order to represent an integrable singularity. In Section V we have defined the Israel–Darmois Israel (1966) junction conditions between a general internal region and a general external BH geometry. In this context, we have shown that satisfying the junction conditions requires the continuity of temperature at the interface, while discontinuities in the tangential pressure signal the presence of phase transitions. It is worth mentioning that these general conditions are also of interest for future studies of other BH exterior regions that possess a central singularity instead of the RN one, for example BHs with cosmic void density profiles Lustosa et al. (2025), BHs surrounded by dark matter Xu et al. (2018), hairy BHs Ovalle et al. (2021), and quintessential BHs Kiselev (2003); Estrada et al. (2025).
Using the framework developed in the previous paragraph, in Sections VI and VII we studied the cases in which the internal geometry of the Reissner–Nordström spacetime is described, respectively, by a string cloud and by our newly proposed string fluid model. In both scenarios, the parameters of the corresponding internal geometries satisfy the conditions listed in Section IV, ensuring the presence of an integrable singularity near the origin, the absence of a potentially unstable de Sitter core, and the absence of an internal horizon that could lead to a breakdown of causality.
In Section VI, we determined the allowed values of the energy density parameter $a$ and of the internal cosmological constant required to satisfy both Israel junction conditions at the interface between the external RN region and the internal string cloud region. In addition, we identified a critical value of the electric charge for which no phase transition occurs in the location of the interface between the two regions.
In Section VII, we derived the conditions that the AdS radius, the electric charge, the screening parameter $b$, and the event horizon must satisfy in order to fulfill both Israel junction conditions at the interface between the external RN region and the internal region described by the new string fluid. Moreover, we numerically determined the critical value of the screening parameter for which no phase transition occurs in the location of the interface between the two regions.

### 16 Mar 2026 v5

@@ Abstract
The Reissner Nordström (RN) black hole is characterized by two well known pathologies: a central singularity and an inner horizon associated with instabilities and a potential loss of predictability. In this work, we show that the RN exterior geometry can arise from an interior spacetime containing an integrable singularity but no inner horizon. In this scenario, tidal forces remain finite near the origin, allowing nondestructive radial infall, while the conventional description in terms of a pointlike mass is replaced by an extended matter distribution. To illustrate this possibility, we provide explicit realizations of such an interior region based on string distributions, namely a cloud of strings (CS) and a newly defined fluid of strings (FS). While the standard cloud of strings model leads to a divergence in the conserved energy associated with timelike Killing vectors, the proposed FS model can be interpreted as a geometrically screened version of the string cloud distribution and admits configurations that, when extended to infinity, describe black holes with finite conserved energy. Physical consistency between the interior region and the RN exterior geometry requires the continuity of temperature across the interface, implying thermal equilibrium between the two regions, while discontinuities in the tangential pressure can signal gravitational phase transitions. These results determine the physical conditions under which string based interior distributions can consistently generate the RN exterior geometry and clarify the circumstances under which phase transitions at the event horizon may arise.
@@ I Introduction
The observation of gravitational waves produced by the merger of two spinning black holes Abbott et al. (2016) has firmly positioned black holes among the most compelling phenomena in gravitational physics. Black hole solutions are typically associated with the presence of a central singularity, where the geometric description of spacetime breaks down due to the divergence of curvature invariants. A widely studied approach to address this issue is provided by regular black holes (RBHs), in which the central singularity is usually replaced by a de Sitter core. However, these models generally possess an inner horizon, whose presence leads to well-known dynamical problems. In particular, the phenomenon of mass inflation can generate severe instabilities at this horizon Poisson and Israel (1990); Brown et al. (2011). More recent studies have reinforced this conclusion, indicating that instabilities associated with the inner horizon appear to be an unavoidable feature of astrophysically relevant RBHs Carballo-Rubio et al. (2023). Moreover, it has been argued that the existence of such an inner horizon may compromise the predictability of spacetime beyond it Ovalle (2023).
An alternative route that has recently attracted attention consists of replacing the regular core with an integrable singularity (IS) Lukash and Strokov (2013). In this type of configuration, curvature invariants may diverge at the origin, while their volume integrals remain finite. An important consequence of this property is that tidal forces experienced by infalling observers can remain finite, suggesting that objects approaching the singularity radially may not necessarily be destroyed. Recent studies, such as Estrada et al. (2024), have shown that these singularities can be weak according to Tipler’s classification Nolan (2000) and may allow the extension of radial geodesics through the origin . Furthermore, possible quantum interpretations of integrable singularities have also been explored Casadio et al. (2023). Additional applications of this framework can be found in Ovalle (2023); Estrada (2025); Arrechea et al. (2025).
On the other hand, string theory suggests that the fundamental constituents of matter may be described as one-dimensional objects known as strings Younesizadeh and Jokar (2022). In an effective gravitational framework, a macroscopic ensemble of such objects can be modeled as a cloud of strings, whose energy–momentum tensor modifies the surrounding spacetime geometry Letelier (1979). This type of source has been widely used to study black hole models surrounded by string distributions Zafar et al. (2025); Javed et al. (2026); Gogoi et al. (2025); Ahmed et al. (2025); Cai et al. (2025). A natural generalization of this model consists of considering a fluid of strings, in which the energy–momentum tensor includes an effective pressure or tension, adopting a structure similar to that of a perfect fluid Letelier (1983). Recently, this framework has been further generalized through the introduction of new equations of state for string fluids Nunes dos Santos (2025). However, when analyzing the conserved energy associated with timelike Killing vectors, it can be observed that the standard energy density of a cloud of strings leads to divergences in the conserved charge defined in Aoki et al. (2021). This feature suggests that the effective description of string distributions should be reconsidered when studying physically well-defined gravitational configurations. Motivated by this observation, in this work we introduce a modified version of the string-cloud energy density within the framework of a fluid of strings. This modification can be interpreted as a geometrical screening of the original distribution and leads to a finite conserved energy.
On the other hand, it is well known that the Schwarzschild black hole corresponds to a vacuum spacetime endowed with a central singularity. Traditionally, this solution is interpreted as being generated by a pointlike mass located at the origin. However, recent works have explored the possibility that the exterior Schwarzschild geometry may instead arise from more general internal configurations, obtained by matching an interior spacetime to the Schwarzschild exterior at the event horizon Ovalle (2024); Maeda (2025). In this type of construction, the horizon can be interpreted as a surface separating two causally disconnected regions of spacetime, allowing the internal structure to be studied without modifying the geometry observed in the exterior region. This perspective naturally raises important physical questions regarding the nature of the matter capable of generating such an interior region. In particular, it is natural to ask whether similar configurations could also give rise to non-vacuum black hole geometries, such as electrically charged solutions, and whether an appropriate internal structure could simultaneously address the presence of a central singularity as well as the instabilities and loss of predictability typically associated with spacetimes containing an inner horizon.
In this context, it is natural to ask whether a similar configuration can be applied to charged solutions, such as the Reissner–Nordström (RN) black hole. This solution describes an electrically charged black hole and is characterized by the simultaneous presence of a central singularity and an inner horizon. The existence of this inner horizon introduces the well-known problems of instability and loss of predictability discussed above. Motivated by these considerations, in this work we investigate a scenario in which the exterior Reissner–Nordström geometry is generated by an extended interior region containing an integrable singularity and lacking an inner horizon.
To construct this scenario, we first analyze the general conditions that an interior geometry with an integrable singularity must satisfy, as well as the Israel–Darmois matching conditions between an interior region and an exterior black hole solution Israel (1966). We further show that these junction conditions admit a natural thermodynamic interpretation: the temperature must remain continuous across the matching surface, while discontinuities in the tangential pressure can be interpreted as signals of phase transitions.
In order to investigate the nature of the matter capable of generating such an interior region for a Reissner–Nordström black hole, we study configurations in which the interior spacetime is described either by a cloud of strings or by the new fluid-of-strings model introduced in this work, while the exterior region corresponds to the Reissner–Nordström solution. In particular, we analyze the physical conditions that lead to the presence of an integrable singularity and the absence of an inner horizon. Finally, we determine the relations between the parameters of the interior string distributions and the exterior Reissner–Nordström geometry for which the matching conditions are satisfied and the temperature remains continuous at the interface. In addition, we identify the regions of the parameter space for which phase transitions at the event horizon may occur or be absent.
@@
II A new representation of a fluid of strings black hole model featuring an integrable singularity and finite conserved energy
@@ III Black Hole models with an Interior Geometry Featuring an Integrable Singularity
We first analyze the physical configuration of generic matter models for both the interior and exterior regions. In the subsequent sections, we focus on the specific case in which the interior region is described by clouds and fluids of strings, while the exterior region is given by the Reissner–Nordström geometry. It is worth mentioning that the general framework presented in this section could also be of physical interest for future studies of other black hole exterior geometries, beyond RN, that possess a central singularity. Examples include black holes with cosmic void density profiles Lustosa et al. (2025), black holes surrounded by dark matter Xu et al. (2018), hairy black holes Ovalle et al. (2021), and quintessential black holes Kiselev (2003); Estrada et al. (2025).
@@ III.2 Description of the exterior spacetime
IV Representation of the INTERIOR REGION AS AN INTEGRABLE SINGULARITY
@@
V Junction Conditions and Their Thermodynamic Implications
In this section, we first analyze the junction conditions between generic matter models describing both the interior and exterior regions. In the subsequent sections, we focus on the specific cases of interest in this work, in which the interior region is described by clouds and fluids of strings, while the exterior region is given by the Reissner–Nordström geometry. It is worth mentioning that the general junction conditions, together with their physical and thermodynamic implications presented in this section, could also be of interest for future studies of other black hole exterior geometries, beyond RN, that possess a central singularity, as in the examples mentioned in Section III Lustosa et al. (2025); Xu et al. (2018); Ovalle et al. (2021); Kiselev (2003); Estrada et al. (2025).
@@ V Junction Conditions and Their Thermodynamic Implications
VI Scenario 1: A Reissner–Nordström Exterior with a Cloud of Strings as the Interior Region with an Integrable Singularity
@@ VI.2 Junction conditions:
VII Scenario 2: A Reissner–Nordström Exterior with a new fluid of Strings as the Interior Region with an Integrable Singularity
@@ VIII Discussion and Summary
In this work, we have investigated an alternative scenario for the internal structure of the Reissner–Nordström (RN) black hole, motivated by the simultaneous presence of a central non-integrable singularity and an inner horizon acting as a Cauchy horizon, which is commonly associated with instabilities and a potential loss of predictability in General Relativity. We have shown that the exterior RN geometry can instead be generated by an extended interior region containing an integrable singularity and lacking an inner horizon, thus replacing the conventional description based on a pointlike mass located at the origin.
We have analyzed the general conditions required to match an interior region with an exterior black hole geometry through the Israel–Darmois junction conditions Israel (1966). We show that thermodynamic consistency across the interface requires the continuity of temperature, suggesting that the two regions must be in thermal equilibrium, while discontinuities in the tangential pressure can signal the presence of gravitational phase transitions. This thermodynamic interpretation provides an additional physical perspective on the matching between the interior and exterior geometries.
To explore the physical nature of the matter capable of sustaining such an interior geometry while generating a Reissner–Nordström exterior geometry, we considered two explicit realizations based on string distributions: a cloud of strings (CS) and a fluid of strings (FS). In the case of the cloud of strings, following the definition of conserved energy given in Ref. Aoki et al. (2021), the analysis of the conserved quantities associated with timelike Killing vectors reveals that the conserved energy diverges. Motivated by this feature, we introduce a new fluid-of-strings (FS) model in which the energy density can be interpreted as a geometrically screened version of the standard string-cloud distribution. This modification leads to configurations that, when extended to infinity, represent black holes with finite conserved energy, in contrast to the divergent conserved energy typically encountered in the string-cloud black hole.
In this way, for both interior regions constructed from the cloud-of-strings and the new fluid-of-strings distributions, we analyze the physical conditions that lead to the presence of an integrable singularity and the absence of an inner horizon. Furthermore, we determine the relations between the parameters of the two interior string distributions and the exterior Reissner–Nordström geometry for which the temperature remains continuous at the interface, thereby implying thermal equilibrium across the matching surface. In addition, we identify the regions of parameter space for which phase transitions at the event horizon may occur or be absent.
It is worth mentioning that the framework developed in this work is not restricted to the Reissner–Nordström geometry. The general conditions obtained for interior geometries containing integrable singularities, together with their thermodynamic implications, may also be relevant for future studies of other black hole exterior spacetimes that exhibit a destructive central singularity and, in some cases, an inner horizon. In such situations, these exterior geometries could potentially be generated by an interior region endowed with an integrable singularity and lacking an inner horizon. Examples of black hole configurations that could be investigated within the framework presented here include black holes with cosmic void density profiles Lustosa et al. (2025), black holes surrounded by dark matter Xu et al. (2018), hairy black holes Ovalle et al. (2021), and quintessential black holes Kiselev (2003); Estrada et al. (2025).